OpenStudy (anonymous):

Solve the following equation: a). y=7x + 4 b). y=6x + 14 y=9x - 6 2x - 9y =30

OpenStudy (anonymous):

Pottermelon! Go man! You are coming to answer my question!

OpenStudy (anonymous):

Sorry. Leaving now. :)

OpenStudy (anonymous):

Not a good answer!

OpenStudy (anonymous):

Wait, do you want me to answer it?

OpenStudy (anonymous):

matthyj is coming to answer my question!

OpenStudy (anonymous):

a). y=7x + 4 y=9x - 6 subsitute 9 x - 6 = 7 x + 4 9x-7x = 4 + 6 2x = 10 x = 5 y = 7 * 2 + 4 y = 18

OpenStudy (anonymous):

pottermelon-thanks for at least check out my question!

OpenStudy (anonymous):

A: Use substitution. 7x + 4 = 9x - 6. Therefore, -2x = -10. x = 5

OpenStudy (anonymous):

Wait how about b?

OpenStudy (anonymous):

Anybody help me solve x & y for b